If a river flows at a rate of 3mph and a boat travels 60 miles upstream and returns in a total time of 9 hours. How fast would the boat travel be traveling in still water?
Answer by scott8148(5891)

let x="rate in still water" ... t=d/r ... 60/(x-3)+60/(x+3)=9 ... 60(x+3)+60(x-3)=9(x+3)(x-3)

120x=9x^2-81 ... 0=9x^2-120x-81 ... 0=3x^2-40x-27

use quadratic equation to find x